### 11 May 2026 v1

# LiBaGS: Lightweight Boundary Gap Synthesis for Targeted Synthetic Data Selection

Abstract
Synthetic data is useful only when the added samples fill missing parts of the training distribution that matter for the downstream task. We introduce LiBaGS, a lightweight, generator-agnostic method for targeted synthetic training data selection. LiBaGS scores candidate synthetic samples by combining decision-boundary proximity, predictive uncertainty, real-data density, and support validity, so that selected samples are both informative and likely to remain on the real data manifold. We then use a boundary-gap allocation rule that targets sparse but realistic decision-boundary neighborhoods, rather than simply adding more data or selecting only the most uncertain candidates. LiBaGS also learns when enough synthetic samples have been added through a marginal-value stopping rule, assigns softer labels near ambiguous boundaries, and uses a diversity objective to avoid redundant near-duplicate selections. Experiments show that LiBaGS improves accuracy over classical oversampling, hard augmentation, uncertainty-based selection, and targeted-generation selection criteria.
1 Introduction
Synthetic data has become a common way to improve supervised learning when real data is scarce, expensive, imbalanced, private, or locally incomplete. Classical oversampling methods such as SMOTE, Borderline-SMOTE, ADASYN, Safe-Level-SMOTE, KMeans-SMOTE, MWMOTE, and Geometric SMOTE generate synthetic samples by interpolating minority examples, emphasizing boundary-adjacent points, or changing the local sampling geometry [7, 14, 16, 6, 30, 4, 10]. Representation-space methods such as manifold oversampling and DeepSMOTE move this idea into learned features, which is especially relevant for image-like data [5, 9]. At the same time, augmentation methods such as mixup, Manifold Mixup, CutMix, Co-Mixup, CMO, and AugMax improve generalization by creating stronger or more diverse training examples [56, 48, 55, 23, 38, 49]. These methods show that synthetic or augmented data can help, but they also highlight a recurring issue: not every synthetic sample is useful.
This issue has become even more pronounced with modern generative models. GANs and diffusion models can produce large candidate pools, and recent work has shown that diffusion-generated data can improve image classification under the right conditions [12, 19, 3]. Targeted-generation methods go further by using uncertainty, classifier feedback, long-tail signals, hard-negative objectives, or curriculum schedules to generate more informative synthetic samples [37, 18, 26, 31, 2, 52, 32, 15, 24]. They show that useful synthetic data is often hard, uncertain, rare, or classifier-informative. But, many methods still treat usefulness primarily as a local score, a generation prompt, or a curriculum signal. LiBaGS instead asks a much more explicit allocation question: where should synthetic samples be placed, and how many should be kept, after accounting for real-data coverage?
A candidate can be rare but irrelevant if it lies far from the decision boundary. A candidate can be uncertain but redundant if the training set already contains many nearby samples. A candidate can look boundary-like but harmful if it is off the real data support. Useful targeted synthetic data should satisfy three conditions at once:
  1. 1.
    it should lie near a decision-boundary neighborhood,
  2. 2.
    occupy a region where real data is locally sparse, and
  3. 3.
    remain on support, i.e. resemble realistic data from the true distribution.
LiBaGS satisfies the above conditions. Let $\mathcal{Z}$ denote the feature space, let $z\in\mathcal{Z}$ denote the feature representation of an input sample, $p(z)$ the real data density, $q(z)$ the synthetic allocation density, $n$ the number of real samples, and $m$ the number of synthetic samples. Let $r(z)$ be a local boundary-gap value that is high only when a candidate is boundary-relevant, uncertain, and support-valid.
We view classification error near the decision boundary as depending on how much useful data exists in a region. This can be approximated by:
$J_{m}(q)=\int_{\mathcal{Z}}\frac{r(z)}{np(z)+mq(z)}dz.$ (1)
Note that $np(z)+mq(z)$ represents the amount of local evidence from real and synthetic samples. Regions that already contain many real samples (large $p(z)$) benefit less from additional synthetic data. In contrast, regions with few real samples (small $p(z)$) but high boundary gap (large $r(z)$) is where synthetic samples can provide the greatest benefit.
Minimizing 1 gives the following boundary-gap allocation rule:
$q^{*}(z)=\frac{1}{m}\left[\sqrt{\frac{r(z)}{\lambda}}-np(z)\right]_{+},$ (2)
where $\lambda$ is a normalization constant automatically determined by the total synthetic allocation constraint. Intuitively, this rule places more synthetic samples in regions that are important for the decision boundary but poorly covered by real data, and places few or no synthetic samples in regions that already have enough real-data coverage.
LiBaGS is deliberately generator-agnostic and does not require retraining a diffusion model or training a new generator. A generator, interpolator, simulator, or augmentation pipeline proposes candidates, which LiBaGS then scores, filters, and diversifies. This is useful because generator quality varies. If a strong generator proposes many valid candidates, LiBaGS selects the most useful ones. If a weak generator proposes many artifacts, the support validity term and adaptive stopping rule prevent the method from blindly adding them. Thus, LiBaGS is a lightweight allocation rule that can be incorporated into existing synthetic data generating pipelines.
We make four contributions. First, we formulate targeted synthetic training data as a boundary-gap allocation problem. Second, we derive a boundary-gap allocation rule that uses boundary proximity, uncertainty, support validity, and real-density coverage. Third, we provide a practical finite-candidate algorithm with adaptive stopping and diversity-aware selection. Fourth, we empirically compare LiBaGS against classical oversampling, hard augmentation, uncertainty-based selection, and targeted-generation selection criteria. Full proofs are in Appendix A.
2 Related Work
Oversampling and interpolation.
SMOTE and its variants are classical baselines for synthetic training data [7, 14, 16, 6, 30, 10, 4]. These methods are often framed around class imbalance, local minority neighborhoods, or interpolation geometry. LiBaGS is motivated by a related intuition that samples near the decision boundary are often especially useful. However, it differs in two important ways: it does not target boundary proximity alone but also discounts regions that are already well covered by real data and it is generator-agnostic and can select from any proposed synthetic candidate pool, rather than being limited to interpolation among minority-class examples.
Augmentation and mixing.
Mixup, Manifold Mixup, CutMix, Co-Mixup, CMO, and AugMax improve generalization by creating similar, patch-level, context-rich, or hard augmented examples [56, 48, 55, 23, 38, 49]. These methods provide strong augmentation baselines, however, LiBaGS is complementary to them since it can select from augmented candidates by asking whether a candidate fills an under-covered decision-boundary neighborhood.
Targeted synthetic generation.
TSynD optimizes latent representations toward high epistemic uncertainty for medical image classification [37]. Feedback-guided synthesis uses classifier feedback while emphasizing support and diversity [18]. GenDataAgent augments data on the fly using signals related to difficult training samples and gradient-update variance [31]. Deliberate Practice improves synthetic-data scaling by dynamically adding challenging examples [2]. DCDM and DisCL control sample difficulty or synthetic-to-real guidance levels [52, 32]. Longtail Guidance and Generate What Matters target rare, hard, or classifier-useful synthetic samples [15, 24]. Boundary-aware latent interpolation, on-manifold adversarial augmentation, targeted diffusion augmentation, and diffusion-based data augmentation are also related [33, 39, 36, 21, 53, 51]. But, LiBaGS contributes a density-aware allocation and stopping rule that can be applied after candidate generation.
Uncertainty, coverage, and active learning.
Active learning selects informative real examples using uncertainty or coverage [42, 41]. Query synthesis and adversarial active sampling methods synthesize or search near decision boundaries rather than only sampling existing points [50, 34]. Training dynamics methods identify hard, ambiguous, or forgotten examples [46, 45]. LiBaGS synthetic selection also includes support filtering, since some generated candidates may fall outside the real data distribution and therefore be unrealistic or unhelpful for training.
3 Method
3.1 Setup and synthetic candidates
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ be a supervised training set and let $h:\mathcal{X}\to\mathcal{Z}$ be a fixed representation map. A candidate generator produces:
$\mathcal{C}=\{(\tilde{x}_{j},\tilde{y}_{j})\}_{j=1}^{M},\quad z_{j}=h(\tilde{x}_{j}).$ (3)
The generator can be a simulator, local interpolation rule, augmentation pipeline, autoencoder, diffusion model, or any class-conditioned sampler [25, 43, 12, 19, 21, 53, 51, 22]. LiBaGS only assumes access to the finite candidate pool. The goal is to select a subset $S\subseteq\mathcal{C}$ for final training.
For each candidate, LiBaGS estimates four quantities. First, a scoring model produces a class-probability vector $\pi(z)$, commonly used for predictive uncertainty, ensembles, and calibrated confidence scores in active learning / uncertainty estimation [42, 11, 29, 13]. The top-two margin is:
$\Delta(z)=\pi_{(1)}(z)-\pi_{(2)}(z),$ (4)
where smaller values indicate proximity to the decision boundary. To emphasize these boundary-adjacent regions, LiBaGS defines the smooth decision boundary neighborhood weight:
$a_{\tau}(z)=\exp\left(-\frac{\Delta(z)^{2}}{2\tau^{2}}\right),$ (5)
where $\tau>0$ controls how broadly points are considered near the decision boundary, with small $\tau$ meaning narrow focus and large $\tau$ meaning broader boundary neighborhood. In practice, $\tau$ is chosen from the empirical distribution of margin values (e.g., using a lower quantile of $\Delta(z)$ over the candidate pool), so that the weighting adapts to the amount of uncertainty in a given task.
The uncertainty score is the predictive entropy:
$u(z)=H(\pi(z)).$ (6)
The support-validity score $b(z)\in[0,1]$ is high for candidates close to the real representation region, following the general idea that synthetic samples should remain near the real data support [18, 54]. The real-density estimate $\hat{p}(z)$ is computed using nearest-neighbor or kernel density estimates [8, 44]. The boundary-gap importance is then:
$r(z)=a_{\tau}(z)u(z)b(z).$ (7)
3.2 Boundary-gap allocation score
The continuous allocation rule in Eq. (1) leads to the following finite candidate score:
$G_{j}=\left[\sqrt{\frac{r(z_{j})}{\lambda}}-n\hat{p}(z_{j})\right]_{+}.$ (8)
Here, $r(z_{j})$ measures how useful candidate $z_{j}$ is for improving the decision boundary, while $\hat{p}(z_{j})$ measures how well that region is already covered by real training data. The score therefore favors candidates that are both boundary-important and underrepresented in the real dataset.
Intuitively, $G_{j}$ becomes large when a candidate lies in a high-uncertainty boundary region that still lacks sufficient real-data coverage, and becomes small when the candidate lies in a region that is already well covered, is far from the decision boundary, or appears off-support and therefore unlikely to provide useful training information.
3.3 Adaptive stopping
A fixed synthetic budget can be brittle, so LiBaGS instead uses an adaptive stopping rule related to diminishing-return selection and model-selection-by-stopping [20, 35, 27]. Intuitively, once a boundary region has already received enough synthetic coverage, adding additional nearby samples should provide progressively smaller gains.
Consider a local region of feature space with boundary importance $r_{j}$. Let $c_{j}$ denote the amount of existing real-data coverage in that region, and let $t_{j}$ denote the number of synthetic samples already allocated there. The marginal improvement from adding one more synthetic sample is modeled as:
$\Delta_{j}(t_{j})=\frac{r_{j}}{c_{j}+t_{j}}-\frac{r_{j}}{c_{j}+t_{j}+1}=\frac{r_{j}}{(c_{j}+t_{j})(c_{j}+t_{j}+1)}.$ (9)
As more synthetic samples are added to the same region, the marginal gain decreases. LiBaGS keeps adding candidates only while the best remaining marginal gain exceeds a threshold, $\eta>0$, which is chosen from the flattening point of the ordered marginal gain curve. Thus, the number of selected synthetic samples $\widehat{m}=|S|$ is learned from the candidate pool rather than fixed in advance.
3.4 Support filtering, soft labels, and diversity
A candidate with high uncertainty but a small support-validity score $b(z)$ is rejected unless it clears the stopping threshold. This is important when the generator is imperfect, since some boundary-like candidates may lie away from the real data support.
For labels, class-conditioned generators provide a proposed class $c(z)$. Near the boundary, hard labels can be overconfident, so we use the class-probability vector $\pi(z)$ to form a soft label:
$\widetilde{y}(z)=\bigl(1-a_{\tau}(z)\bigr)e_{c(z)}+a_{\tau}(z)\pi(z),$ (10)
where $e_{c(z)}$ is the one-hot generator label and $\pi(z)$ is the class-probability vector from the scoring model. Further from the boundary, $a_{\tau}(z)$ is small, so the generator label is trusted more. Near the boundary, $a_{\tau}(z)$ is large, so the probability vector contributes more, reducing overconfidence in ambiguous regions.
Let $v_{j}=G_{j}b(z_{j})$ and let $k(u,j)$ denote a similarity measure between candidates $u$ and $j$ in feature space, such as a Gaussian kernel or cosine similarity. For a selected set $S\subseteq\mathcal{C}$, LiBaGS maximizes:
$F(S)=\sum_{u\in\mathcal{C}}v_{u}\max_{j\in S}k(u,j).$ (11)
This objective rewards a selected set that covers many high-value candidates while avoiding redundant near-duplicates.
Algorithm 1 Lightweight Boundary Gap Synthesis (LiBaGS)
0: real training set $\mathcal{D}$, candidate generator, encoder $h$, scoring model
0: final classifier, selected synthetic set $S$, and learned count $\widehat{m}$
1: Train the scoring model on the real data only.
2: Generate a candidate pool $\mathcal{C}$ and map candidates to $z=h(x)$.
3: For each candidate, estimate $\pi(z)$, $a_{\tau}(z)$, $u(z)$, $\hat{p}(z)$, and $b(z)$.
4: Compute the boundary-gap allocation score $G_{j}$ using Eq. (8).
5: Compute candidate values $v_{j}=G_{j}b(z_{j})$.
6: Set the stopping threshold $\eta$ using early stopping on marginal-gain curve.
7: Select candidates greedily using Eq. (11) and stop when the best marginal gain is below $\eta$.
8: Assign soft labels to selected candidates using Eq. (10).
9: Train the final classifier on $\mathcal{D}\cup S$.
10:return final classifier, selected synthetic set $S$, and $\widehat{m}=|S|$.
4 Theory
This section explains the main choices in Algorithm 1. The main idea is that useful synthetic candidates should be near the decision boundary, uncertain, supported by real data, diverse, and not already well covered by the real training set.
Theorem 1 (Local risk and boundary-gap allocation).
Assume that, in a small region around $z$, the local sample count is proportional to $np(z)+mq(z)$, where $p$ is the real density and $q$ is the selected synthetic density. If the local error decreases with this count and is weighted by this: $r(z)=a_{\tau}(z)u(z)b(z)$, then the allocation objective is:
$J_{m}(q)=\int_{\mathcal{Z}}\frac{r(z)}{np(z)+mq(z)}\,dz.$ (12)
For fixed $m>0$, for some $\lambda>0$, minimizing $J_{m}(q)$ over $q\geq 0$ with $\int q(z)\,dz=1$ gives:
$q^{*}(z)=\frac{1}{m}\left[\sqrt{\frac{r(z)}{\lambda}}-np(z)\right]_{+}.$ (13)
Proof sketch.
Regions with larger $np(z)+mq(z)$ have lower local error, while regions with larger $r(z)$ are more important for boundary learning. The objective is convex in $q$ and its optimal condition assigns synthetic mass until selected regions have equal marginal benefit. Subtracting the existing real coverage and clipping at zero gives Eq. (13); $G_{j}$ in Eq. (8) is the finite-candidate version.
Theorem 2 (Adaptive stopping).
For a small region $j$, let $c_{j}>0$ be real coverage, $t_{j}$ the number of selected synthetic samples, and $r_{j}\geq 0$ the boundary importance. The gain of adding a sample is:
$\Delta_{j}(t_{j})=\frac{r_{j}}{c_{j}+t_{j}}-\frac{r_{j}}{c_{j}+t_{j}+1}=\frac{r_{j}}{(c_{j}+t_{j})(c_{j}+t_{j}+1)}.$ (14)
Greedy selection with threshold $\eta$ accepts gains $\eta$ and stops when best remaining gain is $\eta$.
Proof sketch.
The gain $\Delta_{j}(t_{j})$ decreases as $t_{j}$ grows, so each region has diminishing returns. $\eta>0$ is chosen from the flattening point of the ordered marginal gain curve. Greedy selection always takes the largest available gain and once that gain is below $\eta$, we stop.
Theorem 3 (Soft-label stability).
Let $\rho(z)=\Pr(Y\,|\,Z=z)$ be the true class distribution. The soft label $\widetilde{y}(z)=\bigl(1-a_{\tau}(z)\bigr)e_{c(z)}+a_{\tau}(z)\pi(z)$ satisfies:
$\|\widetilde{y}(z)-\rho(z)\|_{1}\leq\bigl(1-a_{\tau}(z)\bigr)\|e_{c(z)}-\rho(z)\|_{1}+a_{\tau}(z)\|\pi(z)-\rho(z)\|_{1}.$ (15)
Proof sketch.
The result follows directly from the triangle inequality applied to $\widetilde{y}(z)$.
Theorem 4 (Diversity and support filtering).
Let $F(S)=\sum_{u\in\mathcal{C}}v_{u}\max_{j\in S}k(u,j)$ and let $v_{u}=G_{u}b(z_{u})$, with $v_{u}\geq 0$ and $k(u,j)\geq 0$. Then, $F$ is monotone submodular.
Also, for a candidate $j$,
$F(S\cup\{j\})-F(S)\leq\sum_{u\in\mathcal{C}}v_{u}k(u,j),$ (16)
so candidates with low support-weighted neighborhood value cannot pass the stopping threshold.
Proof sketch.
Adding a candidate can only increase coverage, so $F$ is monotone. Adding a candidate also has diminishing returns, so $F$ is submodular. Since $v_{u}=G_{u}b(z_{u})$, low-support regions have small value and are filtered unless they still clear the marginal gain threshold.
5 Experiments
We evaluate LiBaGS in three settings: a two moons boundary-gap task, an $\times8$ handwritten-digit 3 vs. 8 task, and a CIFAR-10 cat vs. dog image task. Across all experiments, LiBaGS uses the same procedure described in Section 3: a scoring model is trained only on the real training data, candidates are scored with $r(z)=a_{\tau}(z)u(z)b(z)$, the allocation score $G_{j}$ is computed, candidates are selected by the objective with marginal-gain stopping, and selected candidates are assigned the soft label in Eq. (10). The selected synthetic count $\widehat{m}$ is learned independently on each seed and is not fixed in advance. For fair comparison, all fixed-count baselines receive the same learned synthetic count $\widehat{m}$ as LiBaGS on a given seed.
Datasets and setup.
The two moons experiment is a nonlinear classification problem, where we draw real training points away from a central boundary region, so the training set contains a local coverage gap near the decision boundary. The candidate pool contains simulator-generated boundary candidates, additional supported candidates, and off-support candidates. The representation map $h$ is a fixed random Fourier feature map, and the final classifier is logistic regression.
The $\times8$ digits experiment uses the scikit-learn handwritten digits data [1], licensed under (CC BY 4.0), restricted to a binary 3 vs. 8 task (since these two numbers may have a difficult decision boundary). The real training set contains a small number of easy examples per class, while the candidate pool contains held-out hard digits, same-class interpolations, and low-structure digits. This setting tests whether LiBaGS can use the boundary-gap score to identify useful ambiguous candidates while avoiding off-support candidates.
For CIFAR-10 [28], which is widely used for research purposes, we use the cat vs. dog task with $25$ real training images per class, $1000$ auxiliary source images per class, and $800$ test images per class. We use a frozen ResNet-50 [17] feature extractor from timm. The candidate pool is generated with Stable Diffusion v1-5 text/image-to-image model [40] using $50$ candidates per class, diffusion strength $0.35$, guidance scale $6.0$, and $30$ denoising steps. The generator is used only to produce the finite candidate pool; LiBaGS then selects a subset adaptively.
Baselines.
We compare against empirical risk minimization (ERM) [47], random candidate selection, noise augmentation, and the same selection-signal baselines used in the CIFAR experiment: TSynD epistemic criterion [37], feedback-guided criterion [18], GenDataAgent gradient-variance criterion [31], Deliberate Practice entropy criterion [2], conformal filtering [54], long-tail rare/hard criterion [15], C2I boundary-influence criterion [24], and uncertainty-only [42]. On the two small datasets, we additionally include classical oversampling and augmentation baselines: SMOTE [7], Borderline-SMOTE [14], ADASYN [16], Safe-Level-SMOTE [6], KMeans-SMOTE [30], DeepSMOTE [9], and AugMax [49]. These additional baselines do not use the external candidate pool, but synthesize examples directly from the real training set.
Table 1: Accuracy across the three datasets across baselines used in the CIFAR-10 experiment. All non-ERM methods use the adaptive synthetic count learned by LiBaGS on each seed; means and standard deviations are over 5 seeds.
Method Two moons 8x8 digits CIFAR-10 cat/dog
ERM [47] 0.8859 $\pm$ 0.0165 0.9391 $\pm$ 0.0167 0.8040 $\pm$ 0.0230
Noise augmentation 0.8977 $\pm$ 0.0226 0.9342 $\pm$ 0.0189 0.7918 $\pm$ 0.0267
Random candidates 0.9299 $\pm$ 0.0198 0.9491 $\pm$ 0.0167 0.8120 $\pm$ 0.0159
TSynD epistemic criterion [37] 0.9021 $\pm$ 0.0313 0.9503 $\pm$ 0.0158 0.8054 $\pm$ 0.0173
Feedback-guided criterion [18] 0.9330 $\pm$ 0.0184 0.9516 $\pm$ 0.0265 0.8078 $\pm$ 0.0158
GenDataAgent gradient-variance criterion [31] 0.9097 $\pm$ 0.0840 0.9342 $\pm$ 0.0306 0.8114 $\pm$ 0.0150
Deliberate Practice entropy criterion [2] 0.9437 $\pm$ 0.0169 0.9466 $\pm$ 0.0303 0.8101 $\pm$ 0.0140
Conformal filtering [54] 0.9209 $\pm$ 0.0254 0.9466 $\pm$ 0.0273 0.8130 $\pm$ 0.0151
Long-tail rare/hard criterion [15] 0.9369 $\pm$ 0.0197 0.9478 $\pm$ 0.0143 0.8088 $\pm$ 0.0167
C2I boundary-influence criterion [24] 0.9429 $\pm$ 0.0170 0.9404 $\pm$ 0.0222 0.8141 $\pm$ 0.0149
Uncertainty-only [42] 0.9431 $\pm$ 0.0166 0.9391 $\pm$ 0.0178 0.8084 $\pm$ 0.0199
LiBaGS 0.9454 $\pm$ 0.0163 0.9540 $\pm$ 0.0315 0.8159 $\pm$ 0.0167
Main results.
Table 1 reports accuracy for the shared baselines across all three datasets. LiBaGS obtains the best mean accuracy in each setting. On two moons, LiBaGS improves over ERM from $0.8859$ to $0.9454$, showing that the boundary-gap allocation score recovers useful samples from the missing boundary region. On $\times8$ digits 3 vs. 8, LiBaGS improves over ERM from $0.9391$ to $0.9540$, while also outperforming the other baselines. On CIFAR-10 cat vs. dog, LiBaGS reaches $0.8159$, compared with $0.8040$ for ERM. These gains are modest but consistent, and they match the intended use case: LiBaGS is most useful when the candidate pool contains boundary-adjacent samples that are informative, on-support, and not already well covered by the real data. Together, the results suggest that explicitly targeting sparse boundary regions can improve synthetic-data selection without requiring a fixed (unknown) synthetic budget.
Tables 2 and 3 give the full results for the two moons experiment and the $\times8$ digits experiment, including the AUROC. The two moons experiment shows the largest absolute improvement because the missing boundary region is explicit. Figure 1 shows this explicitly. Although both tasks have several methods obtaining very similar ranking performance, LiBaGS is still the strongest on accuracy, while some uncertainty baselines have slightly higher AUROC. This suggests that LiBaGS is most useful as a targeted training-data selector: rather than simply improving global ranking of synthetic samples, it adds training signal in sparse boundary regions where additional examples can directly improve the learned classifier.
Table 2: Two moons results. LiBaGS learns the synthetic count (177.2 $\pm$ 42.5 across 5 seeds) by marginal-gain stopping; all baselines use this count for fair selection comparison.
Method Accuracy AUROC
ERM [47] 0.8859 $\pm$ 0.0165 0.9554 $\pm$ 0.0124
Noise augmentation 0.8977 $\pm$ 0.0226 0.9626 $\pm$ 0.0149
Random candidates 0.9299 $\pm$ 0.0198 0.9800 $\pm$ 0.0102
SMOTE [7] 0.8960 $\pm$ 0.0197 0.9604 $\pm$ 0.0152
Borderline-SMOTE [14] 0.9117 $\pm$ 0.0212 0.9671 $\pm$ 0.0145
ADASYN [16] 0.9083 $\pm$ 0.0172 0.9652 $\pm$ 0.0145
Safe-Level-SMOTE [6] 0.8939 $\pm$ 0.0145 0.9584 $\pm$ 0.0101
KMeans-SMOTE [30] 0.8957 $\pm$ 0.0182 0.9598 $\pm$ 0.0143
DeepSMOTE [9] 0.8967 $\pm$ 0.0202 0.9601 $\pm$ 0.0144
AugMax [49] 0.9086 $\pm$ 0.0198 0.9660 $\pm$ 0.0153
TSynD epistemic criterion [37] 0.9021 $\pm$ 0.0313 0.9615 $\pm$ 0.0239
Feedback-guided criterion [18] 0.9330 $\pm$ 0.0184 0.9799 $\pm$ 0.0137
GenDataAgent gradient-variance criterion [31] 0.9097 $\pm$ 0.0840 0.9666 $\pm$ 0.0475
Deliberate Practice entropy criterion [2] 0.9437 $\pm$ 0.0169 0.9857 $\pm$ 0.0083
Conformal filtering [54] 0.9209 $\pm$ 0.0254 0.9725 $\pm$ 0.0194
Long-tail rare/hard criterion [15] 0.9369 $\pm$ 0.0197 0.9831 $\pm$ 0.0112
C2I boundary-influence criterion [24] 0.9429 $\pm$ 0.0170 0.9855 $\pm$ 0.0080
Uncertainty-only [42] 0.9431 $\pm$ 0.0166 0.9855 $\pm$ 0.0083
LiBaGS (ours) 0.9454 $\pm$ 0.0163 0.9853 $\pm$ 0.0082
Table 3: $\times8$ digits results on classes 3 vs. 8. LiBaGS learns the synthetic count (47.4 $\pm$ 10.2 across 5 seeds) by marginal-gain stopping; all baselines use this count for fair selection comparison.
Method Accuracy AUROC
ERM [47] 0.9391 $\pm$ 0.0167 0.9905 $\pm$ 0.0049
Noise augmentation 0.9342 $\pm$ 0.0189 0.9896 $\pm$ 0.0056
Random candidates 0.9491 $\pm$ 0.0167 0.9845 $\pm$ 0.0075
SMOTE [7] 0.9441 $\pm$ 0.0237 0.9915 $\pm$ 0.0059
Borderline-SMOTE [14] 0.9404 $\pm$ 0.0227 0.9895 $\pm$ 0.0065
ADASYN [16] 0.9478 $\pm$ 0.0209 0.9905 $\pm$ 0.0053
Safe-Level-SMOTE [6] 0.9453 $\pm$ 0.0221 0.9896 $\pm$ 0.0067
KMeans-SMOTE [30] 0.9453 $\pm$ 0.0161 0.9908 $\pm$ 0.0055
DeepSMOTE [9] 0.9366 $\pm$ 0.0217 0.9893 $\pm$ 0.0066
AugMax [49] 0.9404 $\pm$ 0.0251 0.9906 $\pm$ 0.0060
TSynD epistemic criterion [37] 0.9503 $\pm$ 0.0158 0.9859 $\pm$ 0.0080
Feedback-guided criterion [18] 0.9516 $\pm$ 0.0265 0.9860 $\pm$ 0.0203
GenDataAgent gradient-variance criterion [31] 0.9342 $\pm$ 0.0306 0.9845 $\pm$ 0.0106
Deliberate Practice entropy criterion [2] 0.9466 $\pm$ 0.0303 0.9911 $\pm$ 0.0072
Conformal filtering [54] 0.9466 $\pm$ 0.0273 0.9856 $\pm$ 0.0248
Long-tail rare/hard criterion [15] 0.9478 $\pm$ 0.0143 0.9914 $\pm$ 0.0063
C2I boundary-influence criterion [24] 0.9404 $\pm$ 0.0222 0.9904 $\pm$ 0.0065
Uncertainty-only [42] 0.9391 $\pm$ 0.0178 0.9917 $\pm$ 0.0051
LiBaGS (ours) 0.9540 $\pm$ 0.0315 0.9873 $\pm$ 0.0104
Figure 1: Qualitative illustration of boundary-gap selection on the two-moons task. Top: with only a small real training set, ERM learns an imperfect decision boundary because several boundary-neighborhood regions are weakly covered by real samples. Bottom: LiBaGS identifies synthetic candidates with high boundary-gap value, combining boundary proximity, uncertainty, support validity, and low real-data coverage. The selected candidates, shown as green and red stars, are added to the real training set and produce a visibly improved decision boundary. The background shading represents the predicted class-probability surface.
6 Computational cost
Let $N$ be the number of real samples, $M$ the number of synthetic candidates, and $d$ the feature dimension. After candidate generation, LiBaGS has three main costs: fitting a lightweight scoring model on real features, scoring estimation for candidates, and diversity and support filtering, with the last step dominating the cost. The diversity and support filtering forms a candidate-candidate similarity matrix, costing $O(M^{2}d)$ time and $O(M^{2})$ memory. Thus, after generation, the exact LiBaGS cost is $O(M^{2}d)$ for the asymptotic time complexity and $O(M^{2})$ for the asymptotic space complexity. Empirically, for image data, generation still takes the most time and space. We used an NVIDIA A100 GPU with 40 GB of VRAM for the CIFAR-10 experiments, including image generation, which took about 1:20 minutes per seed run, given our run settings.
7 Discussion & Impact
LiBaGS can reduce the amount of real data needed in low-resource settings by selecting synthetic examples that target sparse boundary regions. This is especially useful when data collection is expensive, privacy-sensitive, time-consuming, or difficult to repeat. Rather than adding synthetic samples uniformly or relying only on uncertainty, LiBaGS attempts to identify candidates that are simultaneously near the decision boundary, plausible under the real data distribution, and located in regions with limited real-data coverage. This makes synthetic augmentation more targeted and can reduce unnecessary training on redundant or low-value generated samples.
The main risks come from synthetic data generation itself. If a generator reflects bias, label noise, spurious correlations, or unrealistic artifacts, LiBaGS may still select candidates that reinforce those problems. Support filtering reduces this risk but does not eliminate it, since a biased or incomplete real dataset can make biased synthetic samples appear on-support. Similarly, soft labels may reduce overconfidence near the boundary, but they cannot correct systematic errors in the generator or in the scoring model. As a result, LiBaGS should be viewed as a selection method rather than a guarantee of data quality.
In sensitive domains such as medicine, autonomous systems, education, etc., synthetic data can create a false sense of coverage. A selected synthetic sample may appear useful according to model-based scores while still being clinically implausible, demographically biased, or inconsistent with real deployment conditions. In such cases, LiBaGS should be paired with domain review, dataset audits, bias evaluation, external validation, and careful documentation of the generator, candidate pool, selected samples, and intended use.
8 Limitations
LiBaGS depends on the quality of the candidate pool. If the generator produces mostly easy examples, redundant examples, or off-support artifacts, the method may learn a small synthetic count or select candidates that provide only marginal gains. The method also relies on the representation map $h$ and the scoring model; poor representations can make $\hat{p}(z)$, $b(z)$, and $\pi(z)$ unreliable. Future work should compare the baselines in large-scale or multi-class regimes. We also use selection-signal baselines rather than reproducing every full external pipeline, which isolates the relevant selection criteria but may not capture all implementation details of those methods in detail.
9 Conclusion
LiBaGS is a lightweight method for targeted synthetic data selection. Rather than adding a fixed number of generated samples, it scores candidates by boundary importance, uncertainty, support validity, and real-data coverage, then stops when the estimated marginal gain becomes small. Across the two moons task, the $\times8$ digits 3 vs. 8 task, and the CIFAR-10 cat vs. dog experiment, LiBaGS improves over ERM and achieves the best mean accuracy among several selection baselines. The results support the claim that synthetic data is most useful when it fills sparse, on-support boundary gaps rather than merely increasing the size of the dataset.
Appendix A Technical appendices and supplementary material
A.1 Proof of Theorem 1
Theorem 1 (Local risk and boundary-gap allocation).
Assume that, in a small region around $z$, the local sample count is proportional to $np(z)+mq(z)$, where $p$ is the real density and $q$ is the selected synthetic density. If the local error decreases with this count and is weighted by this: $r(z)=a_{\tau}(z)u(z)b(z)$, then the allocation objective is:
$J_{m}(q)=\int_{\mathcal{Z}}\frac{r(z)}{np(z)+mq(z)}\,dz.$ (17)
For fixed $m>0$, for some $\lambda>0$, minimizing $J_{m}(q)$ over $q\geq 0$ with $\int q(z)\,dz=1$ gives:
$q^{*}(z)=\frac{1}{m}\left[\sqrt{\frac{r(z)}{\lambda}}-np(z)\right]_{+}.$ (18)
Proof.
In a small region around $z$, the number of nearby training examples is proportional to $np(z)+mq(z)$. Local class-probability estimates become more stable as this count increases, so the local error is proportional to the inverse count. LiBaGS weights this local term using $r(z)=a_{\tau}(z)u(z)b(z)$, where $a_{\tau}(z)$ emphasizes boundary regions, $u(z)$ measures uncertainty, and $b(z)$ measures support. This gives the objective:
$J_{m}(q)=\int_{\mathcal{Z}}\frac{r(z)}{np(z)+mq(z)}\,dz.$ (19)
We optimize over valid synthetic allocation densities. The density $q(z)$ specifies where the fixed synthetic budget $m$ is placed in representation space. Thus $q(z)\geq 0$ and $\int q(z)\,dz=1$, while $mq(z)$ is the amount of synthetic mass assigned near $z$.
For a fixed $z$, define $\phi_{z}(q)=\frac{r(z)}{np(z)+mq}$. Since $\phi_{z}^{\prime\prime}(q)=\frac{2m^{2}r(z)}{(np(z)+mq)^{3}}\geq 0$, each local term is convex in the allocation variable. Therefore, the full objective is convex in $q$.
Let $q^{*}$ be the optimal allocation. A region with $q^{*}(z)>0$ receives synthetic mass. Since the total synthetic budget is fixed, moving a small amount of mass between any two such regions should not reduce the objective; otherwise, $q^{*}$ would not be optimal. The derivative $\frac{d}{dq(z)}\phi_{z}(q)$ measures local change from adding synthetic mass near $z$. It is negative because adding mass increases the denominator $np(z)+mq(z)$ and decreases the local objective term. If two selected regions had different derivatives, shifting mass toward the region with the larger decrease would improve allocation. So, all regions with $q^{*}(z)>0$ must have the same derivative value. Hence, for some constant $C>0$,
$\frac{d}{dq(z)}\left(\frac{r(z)}{np(z)+mq(z)}\right)=-C.$ (20)
Differentiating gives:
$\frac{mr(z)}{(np(z)+mq^{*}(z))^{2}}+C=0.$ (21)
Solving for the total local coverage gives:
$np(z)+mq^{*}(z)=\sqrt{\frac{mr(z)}{C}}.$ (22)
Absorbing constants into $\lambda=\frac{C}{m}>0$ yields:
$mq^{*}(z)=\sqrt{\frac{r(z)}{\lambda}}-np(z).$ (23)
Since synthetic density cannot be negative, negative allocations are clipped to zero:
$q^{*}(z)=\frac{1}{m}\left[\sqrt{\frac{r(z)}{\lambda}}-np(z)\right]_{+},$ (24)
where $\lambda>0$ is chosen so that the allocation satisfies $\int q^{*}(z)\,dz=1$.
Replacing $z$ with $z_{j}$ and replacing $p(z_{j})$ with the empirical density estimate $\hat{p}(z_{j})$ gives the finite-candidate score in Eq. (8). ∎
A.2 Proof of Theorem 2
Theorem 2 (Adaptive stopping).
For a small region $j$, let $c_{j}>0$ be real coverage, $t_{j}$ the number of selected synthetic samples, and $r_{j}\geq 0$ the boundary importance. The gain from adding one more sample to region $j$ is:
$\Delta_{j}(t_{j})=\frac{r_{j}}{c_{j}+t_{j}}-\frac{r_{j}}{c_{j}+t_{j}+1}=\frac{r_{j}}{(c_{j}+t_{j})(c_{j}+t_{j}+1)}.$ (25)
Greedy selection with threshold $\eta$ accepts gains $\eta$ and stops when best remaining gain is $\eta$.
Proof.
For region $j$, adding one synthetic sample changes the local term from $\frac{r_{j}}{c_{j}+t_{j}}$ to $\frac{r_{j}}{c_{j}+t_{j}+1}$. The gain is therefore:
$\Delta_{j}(t_{j})=\frac{r_{j}}{c_{j}+t_{j}}-\frac{r_{j}}{c_{j}+t_{j}+1}=\frac{r_{j}}{(c_{j}+t_{j})(c_{j}+t_{j}+1)}.$ (26)
As $t_{j}$ increases, the denominator increases, so the marginal gain decreases. Thus, each region has diminishing returns: the first synthetic samples assigned to an under-covered region are more useful than later ones assigned to the same region.
With some threshold $\eta>0$ (where $\eta$ is chosen from the flattening point of the ordered marginal gain curve), adding a sample is worthwhile only if its gain is above $\eta$. Greedy selection always chooses the largest available gain across all regions. If this largest remaining gain is below $\eta$, then every other remaining gain is also below $\eta$, so no additional synthetic sample clears the stopping criterion.
Therefore, the greedy rule accepts exactly the samples whose estimated marginal gain is above $\eta$ and stops once all remaining gains are too small. This proves the adaptive stopping rule and shows why the selected count $\widehat{m}$ is determined by the data rather than fixed in advance. ∎
A.3 Proof of Theorem 3
Theorem 3 (Soft-label stability).
Let $\rho(z)=\Pr(Y\,|\,Z=z)$ be the true class distribution. The soft label $\widetilde{y}(z)=\bigl(1-a_{\tau}(z)\bigr)e_{c(z)}+a_{\tau}(z)\pi(z)$ satisfies:
$\|\widetilde{y}(z)-\rho(z)\|_{1}\leq\bigl(1-a_{\tau}(z)\bigr)\|e_{c(z)}-\rho(z)\|_{1}+a_{\tau}(z)\|\pi(z)-\rho(z)\|_{1}.$ (27)
Proof.
The soft label is:
$\widetilde{y}(z)=(1-a_{\tau}(z))e_{c(z)}+a_{\tau}(z)\pi(z).$ (28)
Subtracting the true class distribution $\rho(z)$ on both sides gives:
$\widetilde{y}(z)-\rho(z)=(1-a_{\tau}(z))(e_{c(z)}-\rho(z))+a_{\tau}(z)(\pi(z)-\rho(z)).$ (29)
Taking the $\ell_{1}$ norm and applying the triangle inequality, we get:
$\|\widetilde{y}(z)-\rho(z)\|_{1}\leq(1-a_{\tau}(z))\|e_{c(z)}-\rho(z)\|_{1}+a_{\tau}(z)\|\pi(z)-\rho(z)\|_{1}.$ (30)
A.4 Proof of Theorem 4
Theorem 4 (Diversity and support filtering).
Let $F(S)=\sum_{u\in\mathcal{C}}v_{u}\max_{j\in S}k(u,j)$ and also let $v_{u}=G_{u}b(z_{u})$, with $v_{u}\geq 0$ and $k(u,j)\geq 0$. Then, $F$ is monotone submodular.
Also, for a candidate $j$,
$F(S\cup\{j\})-F(S)\leq\sum_{u\in\mathcal{C}}v_{u}k(u,j),$ (31)
so candidates with low support-weighted neighborhood value cannot pass the stopping threshold.
Proof.
Let $A\subseteq B$, then for every $u$,
$\max_{j\in A}k(u,j)\leq\max_{j\in B}k(u,j).$ (32)
Since $v_{u}\geq 0$, summing over $u$ gives $F(A)\leq F(B)$. Thus, $F$ is monotone.
Let $A\subseteq B$ and $e\notin A,B$. For every $u$, let $m_{A}(u)=\max_{j\in A}k(u,j),m_{B}(u)=\max_{j\in B}k(u,j)$.
Since $A\subseteq B$, we have $m_{A}(u)\leq m_{B}(u)$ for every $u$. The marginal gain of adding $i$ to $A$ is:
$F(A\cup\{i\})-F(A)=\sum_{u\in\mathcal{U}}v_{u}\left[\max\{m_{A}(u),k(u,i)\}-m_{A}(u)\right].$
Equivalently,
$F(A\cup\{i\})-F(A)=\sum_{u\in\mathcal{U}}v_{u}\left[k(u,i)-m_{A}(u)\right]_{+},$
where $x]_{+}$ clips negative values to 0. Similarly,
$F(B\cup\{i\})-F(B)=\sum_{u\in\mathcal{U}}v_{u}\left[k(u,i)-m_{B}(u)\right]_{+}.$
Because $m_{A}(u)\leq m_{B}(u)$, we have:
$\left[k(u,i)-m_{A}(u)\right]_{+}\geq\left[k(u,i)-m_{B}(u)\right]_{+}$
for every $u$. Since $v_{u}\geq 0$, summing over $u$ gives:
$F(A\cup\{i\})-F(A)\geq F(B\cup\{i\})-F(B).$
Therefore, $F$ is submodular.
So, $F$ is monotone submodular, i.e. adding candidates always improves coverage, but each additional candidate gives increasingly smaller gains once similar candidates have already been selected.
Finally, the marginal gain of adding $j$, as also shown above, is:
$F(S\cup\{j\})-F(S)=\sum_{u\in\mathcal{C}}v_{u}\left[k(u,j)-\max_{\ell\in S}k(u,\ell)\right]_{+}.$ (33)
Since $x-y]_{+}\leq x$ for $x\geq 0$,
$F(S\cup\{j\})-F(S)\leq\sum_{u\in\mathcal{C}}v_{u}k(u,j).$ (34)
Thus, adding candidate $j$ yields only a small gain when it is highly similar to candidates already contained in the selected set $S$. ∎
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction reflect the proposed LiBaGS method, its theoretical motivation, and the empirical findings presented throughout Section˜1-5.
  5. Guidelines:
    • The answer [N/A] means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A [No] or [N/A] answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses limitations related to candidate pool quality, generator strength, and support estimation in Section˜8.
  10. Guidelines:
    • The answer [N/A] means that the paper has no limitation while the answer [No] means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate “Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theoretical assumptions are stated explicitly in Sections 3 and 4, with proof sketches in the main paper and complete proofs provided in Appendix A.
  15. Guidelines:
    • The answer [N/A] means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 5 provides sufficient implementation, dataset, and training details to reproduce the main experimental results.
  20. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • If the paper includes experiments, a [No] answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: All of the data is publicly available and the supplementary material includes LiBaGS-specific code. We will release all the code upon acceptance.
  25. Guidelines:
    • The answer [N/A] means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so [No] is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specifies the datasets, architectures, generators, and evaluation in Section 5.
  30. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All of the experiments in Section˜5 report the mean and standard deviation of the accuracy and AUROC across 5 runs with different seeds.
  35. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The authors should answer [Yes] if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, the authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Compute resource details are provided in Section˜6.
  40. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer [N/A] means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer [No] , they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Section 7 discusses both the potential benefits of improved synthetic data selection and possible risks such as bias amplification or misuse of generative models.
  50. Guidelines:
    • The answer [N/A] means that there is no societal impact of the work performed.
    • If the authors answer [N/A] or [No] , they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate Deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release any generative models. Our method is generator-agnostic and can be used in conjunction with a custom generator or a pre-trained generator to synthesize data.
  55. Guidelines:
    • The answer [N/A] means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets, pretrained models, and external assets are properly cited in Section˜5.
  60. Guidelines:
    • The answer [N/A] means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide LiBaGS-specific code in the supplementary material.
  65. Guidelines:
    • The answer [N/A] means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper uses existing datasets and does not involve crowdsourcing nor new research with human subjects.
  70. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This research does not involve crowdsourcing nor new research with human subjects.
  75. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigor, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer [N/A] means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy in the NeurIPS handbook for what should or should not be described.

### 13 May 2026 v2

@@
 Thus, adding candidate $j$ yields only a small gain when it is highly similar to candidates already contained in the selected set $S$. ∎
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The abstract and introduction reflect the proposed LiBaGS method, its theoretical motivation, and the empirical findings presented throughout Section˜1-5.
-  5. Guidelines:
-    • The answer [N/A] means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A [No] or [N/A] answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: The paper discusses limitations related to candidate pool quality, generator strength, and support estimation in Section˜8.
-  10. Guidelines:
-    • The answer [N/A] means that the paper has no limitation while the answer [No] means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate “Limitations” section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: All theoretical assumptions are stated explicitly in Sections 3 and 4, with proof sketches in the main paper and complete proofs provided in Appendix A.
-  15. Guidelines:
-    • The answer [N/A] means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Section 5 provides sufficient implementation, dataset, and training details to reproduce the main experimental results.
-  20. Guidelines:
-    • The answer [N/A] means that the paper does not include experiments.
-    • If the paper includes experiments, a [No] answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24. Justification: All of the data is publicly available and the supplementary material includes LiBaGS-specific code. We will release all the code upon acceptance.
-  25. Guidelines:
-    • The answer [N/A] means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so [No] is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The paper specifies the datasets, architectures, generators, and evaluation in Section 5.
-  30. Guidelines:
-    • The answer [N/A] means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: All of the experiments in Section˜5 report the mean and standard deviation of the accuracy and AUROC across 5 runs with different seeds.
-  35. Guidelines:
-    • The answer [N/A] means that the paper does not include experiments.
-    • The authors should answer [Yes] if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
-    • If error bars are reported in tables or plots, the authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Compute resource details are provided in Section˜6.
-  40. Guidelines:
-    • The answer [N/A] means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research conforms to the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer [N/A] means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer [No] , they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: Section 7 discusses both the potential benefits of improved synthetic data selection and possible risks such as bias amplification or misuse of generative models.
-  50. Guidelines:
-    • The answer [N/A] means that there is no societal impact of the work performed.
-    • If the authors answer [N/A] or [No] , they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate Deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: We do not release any generative models. Our method is generator-agnostic and can be used in conjunction with a custom generator or a pre-trained generator to synthesize data.
-  55. Guidelines:
-    • The answer [N/A] means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: All datasets, pretrained models, and external assets are properly cited in Section˜5.
-  60. Guidelines:
-    • The answer [N/A] means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We provide LiBaGS-specific code in the supplementary material.
-  65. Guidelines:
-    • The answer [N/A] means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper uses existing datasets and does not involve crowdsourcing nor new research with human subjects.
-  70. Guidelines:
-    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: This research does not involve crowdsourcing nor new research with human subjects.
-  75. Guidelines:
-    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigor, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: This research does not involve LLMs as any important, original, or non-standard components.
-  80. Guidelines:
-    • The answer [N/A] means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy in the NeurIPS handbook for what should or should not be described.